Continued to work on An/Ar.13. We have not had a soil change, but we continue to have material. There maybe a wall in the center of the trench, but at this point it is still difficult to distinguish between rock fall. We continue to dig in order to see if their are stacked rocks. Drew soil profile for Ar/An.Op4b.